OpenStudy (anonymous):

PLEASE!!!! I know I have the formula done correctly but I'm not coming up with the right answer: what is the sum of the geometric series. Sigma 6(2)^x where n=1 and 10 in the series. I have 6(2)^x +6(2)^x....10 times. Answer choices are a:15,658 b: 6,138 c:12,276 d: 756 I keep coming up with 1,440

OpenStudy (anonymous):

\[\sum_{n=1}^{10}6(2)^x\]

OpenStudy (mrnood):

I think this is written confusingly. Normally it would be x=1 > x=10. i.e the symbol under the sigma sign should read x=1. (alternatively it should be 2^n rather than 2^x). If that is the case then it means : for each value of x from 1 to 10 add up the value of 6(2)^x Note that as the 6 is constant you can remove it from the Summation and 'multiply' so it becomes 6* sum(2^1,2^2...........2^10) You might notice that summing the powers of 2 is like a binary number so you have 11111111110 in binary (the 0 is because there is no 2^0 term) convert and multiply by 6....

OpenStudy (anonymous):

I agree it is written strangely. And that they should all be either "x" or "n". But you lost me with the explanation. I really have no idea how to get to one of the answers listed.

OpenStudy (anonymous):

so it would be: 6(2)+6(4)+6(8)....out 10 terms?

OpenStudy (mrnood):

Yes that is correct - and it does come to one of the answers. My binary explanation may be a little off the topic - but it does give a quick way to come to the answer. But your way will work too..

OpenStudy (anonymous):

That does work out to one of the answers

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Here a=2, r=2>1,n=10 ∑_(x=1)^10▒〖6(2)^x=6{∑_(x=1)^10▒2^x } 〗=6{2^1+2^2+2^3…+2^10 } Sn=a(r^n-1)/(r-1),n>1 S10=2(2^10-1)/(2-1)=2(2^10-1) Total solution=6*2(2^10-1)=12(1024-1)=12*1023=?